Find the extreme values of the function f(x,y) = 2x^2 + 3y^2 - 4x - 5

on the set D = {(x,y) : x^2 + y^2 < or = 16 }


1
Expert's answer
2021-10-13T13:24:49-0400

Solution

Critical Points and Relative Extrema:

∂f/∂x = 4x – 4 = 0 =>  x = 1

∂f/∂y = 6y = 0 =>  y = 0

fxx = 4 > 0 and fyy = 6 > 0, fxy = 0  => J = fxxfyy - fxy2 >0  => point x = 1, y = 0 is relative minimum

Another extreme values are on the bound of D – circle x2 + y2 = 16

Substituting from equation of circle  y2 = 16 - x2 into the function f(x,y):

f(x) = 2x2 + 3(16 – x2) - 4x – 5  =>  f(x) = 43 – 4x – x2 

f’(x) = -4-2x = 0 => x = -2 => "y\ =\ \pm\sqrt{12}=\pm2\sqrt3"  

So function has a minimum at (x,y) = (1,0) where f(x,y) = -7 and two maximums at "(x,y)\ =\ (-2,\ 2\sqrt3)"  where f(x,y) = 47 and at "(x,y)\ =\ (-2,\ -2\sqrt3)"  where f(x,y) = 47 too.